Circular No. 8589

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2005dm IN IC 219 Further to IAUC 8584, M. Baek and W. Li report the LOSS discov-

ery of an apparent supernova on unfiltered KAIT images taken on Aug. 26.50 (mag 17.6) and 27.51 UT (mag 17.5). SN 2005dm is located at α = 2^{h}18^{m}39^{s}.25, δ = −6^{o}54′10′′.8 (equinox 2000.0), which is 7′′.0 east and 1′′.5 north of the nucleus of IC 219. A KAIT image taken on Feb. 2.16 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2005dn IN NGC 6861F Further to IAUC 8549, C. Jacques reports the discovery by C. Cole-

santi, E. Pimentel, T. Napoleao, and himself of an apparent supernova (mag 15.0) on unfiltered images taken on Aug. 27.03 and 28.95 UT with a 0.30-m f/3 Schmidt-Cassegrain telescope at Belo Horizonte, Brazil, in the course of the survey by the Brazilian Supernovae Search Team. The new object is located at α = 20^{h}11^{m}11^{s}.73, δ = −48^{o}16′35′′.5 (equinox 2000.0), which is 0′′.2 east and 5′′.6 south of the center of NGC 6861F (Sulentic and Tifft 1973, Revised NGC of Nonstellar Astron. Objects). Nothing is visible at this location on a CCD image taken on Apr. 5.23 (limiting mag 19.0) or on a red Palomar Sky Survey image from 1975.

2005 OE_{3}

L. A. M. Benner, J. D. Giorgini, S. J. Ostro, and R. F. Jurgens, Jet Propulsion Laboratory; and M. W. Busch, California Institute of Technology, report: “Goldstone 8560-MHz (3.5-cm) radar observations of 2005 OE_{3} (cf. MPEC 2005-P03) during Aug. 21.05200–21.35451 UTC reveal echoes with a visible range depth of 0.5 μs (75 m) that are unresolved in Doppler frequency at a resolution of 0.025 Hz. If we assume that the true-range extent is double the visible-range extent, then these results establish that P /cos(δ) > 150, where δ is the subradar latitude and P is the rotation period in hours. If the radar view was at least 10^{o} from the pole, then the upper bound on the bandwidth and a diameter of 150 m imply a rotation period of at least 100 hr. Our radar astrometry collapsed the instantaneous 3σ time-delay uncertainty on Aug. 21.00000 from 0.1 s (15000 km) to 10^{−6} s (0.15 km), reduced the 3σ Doppler frequency uncertainty from 1500 to 0.54 Hz, and expanded the interval of reliable close-earth-approach predictions from the current year to the interval from 1462 to 2440.”

2005 August 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green